# WildGraph: Realistic Graph-based Trajectory Generation for Wildlife

### v1

@@ Abstract.
Trajectory generation is an important task in movement studies; it circumvents the privacy, ethical, and technical challenges of collecting real trajectories from the target population. In particular, real trajectories in the wildlife domain are scarce as a result of ethical and environmental constraints of the collection process. In this paper, we consider the problem of generating long-horizon trajectories, akin to wildlife migration, based on a small set of real samples. We propose a hierarchical approach to learn the global movement characteristics of the real dataset, and recursively refine localized regions. Our solution, WildGraph, discretizes the geographic path into a prototype network of H3111https://www.uber.com/blog/h3/ regions and leverages a recurrent variational auto-encoder to probabilistically generate paths over the regions, based on occupancy. WildGraph successfully generates realistic months long trajectories using a sample size as small as 60. Experiments performed on two wildlife migration datasets demonstrate that our proposed method improves the generalization of the generated trajectories in comparison to existing work, while achieving superior or comparable performance in several benchmark metrics. Our code is published on the following repository: https://github.com/aliwister/wildgraph.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Trajectory data provide invaluable insights into the patterns and behaviors of movement, enabling a deeper understanding of the underlying mechanisms driving movement across various domains, including human, vehicle, and wildlife. In particular, wildlife trajectories are utilized by scientists across different fields to promote wildlife conservation (Wall et al., 2014), anti-poaching efforts (Park et al., 2015), climate change studies (McRae et al., 2008), human-wildlife conflict (Buchholtz et al., 2020), and to help guide policy towards wildlife preservation (Tomkiewicz et al., 2010). For example, trajectories help identify where animal movement may intersect with known poaching activity, and hence promote anti-poaching efforts. Similarly shifting in trajectory
However, the collection process of real trajectory data may be associated with various challenges. Wildlife researchers have cited ethical and technical constraints involved in wildlife trajectory data collection (Dore et al., 2020). In particular, data is traditionally collected using devices attached directly to the subject, which requires capturing and possibly sedating it. The devices are further constrained by the size of the animal, and may be prone to intermittent or permanent failures, thus limiting the availability of trajectory data.
@@
In order to overcome challenges associated with collecting real movement data, trajectory generation has been considered in various domains, including human mobility (Long et al., 2023; Yuan et al., 2022; Rossi et al., 2021b), vehicle mobility (Cao and Li, 2021; Rossi et al., 2021a), as well as wildlife (Al-Lawati et al., 2023; Technitis et al., 2015) movement studies. Ideally, we want the generated set of trajectories to have the same characteristics (measured by, e.g. Hausdorff Distance or correlation coefficient) of a set of real trajectories, and generates original samples that preserve the patterns of the real trajectories (e.g., clusters or distance similarity). For example, in Figure 1, it can be visually verified that WildGraph generates much better trajectories, than using a Generative Adversarial Network (GAN) for this problem. The generated trajectories provide valuable insight in experimental studies, scenario planning, and risk assessment, where researchers can simulate the effects of polices and environmental decisions, and thereby improve their robustness and generalization (Beach and Clark, 2015; Catano et al., 2015).
However, unlike human and vehicle trajectories, wildlife trajectory generation presents a unique set of challenges. First is the small data problem: real trajectory datasets are sparse and include a small number of subjects. Second, the length of each trajectory tend to span several months, or years (e.g. in that long timespan, they may vary due to confounding factors, weather patterns, extreme events that are outliers, etc.) to provide practical value. Finally, often there are no established network information, e.g. roads, which wildlife may follow, (although in some land animals there may be tracks on the ground trailed by the migrations, these are not rigid and animals often stray from them to create new trails/paths easily) thus increasing the complexity of the generation task.
Despite these challenges, interestingly, simple deep learning based techniques, such as Variational Auto Encoders (VAEs) have demonstrated promising results in generating long-horizon trajectories using a small set of real training samples (Chen et al., 2021; Al-Lawati et al., 2023). VAEs accomplish this by mapping trajectories in a probabilistic manner to a latent space, and teaching the decoder to reconstruct the encoded trajectory from the latent representation. Hence, new trajectories that have similar properties as the real trajectory may be generated by sampling the latent space.
@@
In this paper, we address these challenges by constructing a prototype network of regions for wildlife movement that models trajectories as a path across the network. We utilize the set of known trajectories to construct the global regions, and recursively refined localized regions, aka nodes, into multiple granular nodes. This helps preserve the global direction of movement, and at the same time captures the local properties of the generated trajectories. By adjusting the region sizes of the network to limit the selection of potential points from a particular region, we improve the generalization of the generated trajectories, and provide a better outcome across multiple benchmarks.
@@
Once each known trajectory is encoded as a path over the network, we process it using a Variational Recurrent Network (VRN), which trains a model that generates a dictionary of latent spaces for each region, as depicted in Figure 3. The details of our framework is discussed in further details in Section 3.
@@
  • We propose WildGraph: a framework that generates realistic long-horizon wildlife trajectories based on a sparse set of real trajectories.
@@
  • We conduct multiple experiments against two wildlife datasets, and benchmark our methods against our earlier work, WildGEN (Al-Lawati et al., 2023), a Variational Auto-encoder (VAE) based approach, Generative Adversarial Network (GAN) based approach, and a Transformers-based approach based on the prototype network.
To the best of our knowledge, this paper along with our previous (unpublished) workshop paper: WildGEN (Al-Lawati et al., 2023) are the first to propose deep-learning based trajectory generation for wildlife movement. Specifically, WildGEN, demonstrated that deep learning approaches yield significantly more realistic trajectories than traditional methods, such as Correlated Random Walk (CRW) (Renshaw and Henderson, 1981) and Levy Walk (Wosniack et al., 2017). This study builds upon our previous work by refining the point selection process along trajectories, as assessed by the similarity between generated trajectories and real trajectories in our test set.
@@ 2. Related Work
In this section, we discuss some recent works that are related to the methods and solutions we present in our paper, from different research areas.
Trajectory Generation. In the pre-deep-learning era, trajectory generation was relatively subject specific. In ecology, researchers used methods such as Levy Walk (Wosniack et al., 2017), Correlated Random Walk (CRW) (Technitis et al., 2015), and other velocity based methods (Technitis et al., 2015; Ahearn et al., 2017), that rely predominantly on the subject (animal) considered. The advent of deep-learning-based generative models such as VAEs and Generative Adversarial Networks (GANs) resulted in some of the earliest attempts at using these approaches for trajectory generation. Chen, et al. (Chen et al., 2021) benchmarked GANs and VAEs for the task of generation, and found that VAEs provide more consistent results. Our earlier work: WildGEN, was among the first methods to address the problem of trajectory generation for wildlife (Al-Lawati et al., 2023). We utilized VAEs with added layers of post-processing that mitigated some of the challenges discussed above. However, much like other VAE-based methods, it suffers from limited generalization capabilities.
@@
Trajectory Similarity. Trajectory similarity is a complementary problem to trajectory generation since the goal of generation is to produce a set of trajectories similar to a set of known trajectories. An interesting approach to the problem was considered by Yao, et al. (Yao et al., 2022), where a region is recursively subdivided to capture the region based similarity between compared trajectories. Similar to our approach, they recursively split regions into smaller regions to create a balanced graph. Next, they utilized a graph-based attention layer to model the embedding of each node. In comparison, our approach utilizes node2vec (Grover and Leskovec, 2016) to overcome the over-parameterization in our problem given the disparity in the training sample sizes. Finally, since the trajectory is considered as a whole in the trajectory similarity problem, a snapshot of the graph may be processed based on the Normalized Lagrangian of the adjacency and degree matrices. A time-aware approach to generate similar graph snapshots at each step in a generation problem would render the problem considerably inefficient. Moreover, it will fail in the presence of cycles in the graph because it cannot determine the correct sequence of nodes in the trajectory.
@@
Graph Generation. Early generative graph models were based on adjacency matrix generation using encoder-decoder structures (You et al., 2018; Simonovsky and Komodakis, 2018). However, these approaches fail to capture the dynamic nature of spatio-temporal graphs. Recently, models such as STGD-VAEs (Du et al., 2022), and STGEN (Ling et al., 2023) were proposed to address these limitation. STGD-VAEs factor graphs into dependent ordinal values and jointly optimize the objective. On the other hand, STGEN utilizes the distribution of spatiotemporal walks to adversarially learn the characteristics of the graph. However, multi-objective approaches appear to be a poor fit for this problem, given the small sample sizes, and long horizon. We validate this in our experiments and demonstrate the GANs provide poor results on all metrics (see  1).
@@
Sequence Modeling. Sequence-to-Sequence models have been used for various tasks in wildlife trajectory studies, including prediction (Feng and Gordon, 2019), imputation (Li et al., 2021) and interpolation (Wan and Dodge, 2023). They identify implicit patterns by using LSTM or RNN memory trained with new observations in an auto-regressive manner (Park et al., 2018). However, the multi-variate nature of spatial data often results in a quick deterioration of predictions on tasks involving long horizon prediction (Sagheer and Kotb, 2019), particularly in small data settings. To address this, Li, et al., (Li et al., 2021) attempted to simplify the problem to a prediction of step size instead of location, but this approach offers very limited usefulness. Our work adopts a more practical approach by leveraging the properties of the H3 spatial library to represent geospatial regions, where each region is encoded with a string representing its address encoded in the library, thus reducing the problem of generating a path to one in a single dimension.
@@
In addition, encoding each region’s connection (i.e., the probability of whether a bird will fly from one region to another) with other regions provides additional context to the problem, and enables us to generate long horizon trajectories based on a small set of training samples.
@@ 3.1. Problem Statement
Given a set of $n$ real trajectories in an area $\mathcal{A}$, each consisting of $m$ points, the objective is to produce a set of synthetic trajectories that maximize the following:
@@
    (Path Similarity) similarity between the points of the synthetic trajectories to that of the real trajectories, and,
@@
Formally, the task is to learn a model to generate synthetic trajectories that maximize the similarity between a set of held-out test-set comprising of $k$ real trajectories, denoted as $T_{1},T_{2},\ldots,T_{k}$, each comprising $m$ spatio-temporal points, where $T_{i}=(p_{1}^{i},p_{2}^{i},\ldots,p_{m}^{i})$, with $p_{j}^{i}$ representing latitude ($y$) and longitude ($x$) values. The objective is to find a corresponding set $S={S_{1},S_{2},\ldots,S_{t}}$ of synthetic trajectories that best matches the (test set of) real trajectories.
To achieve this, we aim to maximize the overall similarity between $T$ and $S$, denoted as:
$\max\text{ Similarity}(T,S)+\max\text{Likeness}(T,S)$
Here, the similarity measure encompasses both Path similarity, and Likeness. Note that using a held-out set to compare with the generated trajectories, post-training, ensures robustness and generalization of the proposed solution. Further details on the evaluation metrics utilized are provided in Section 3.5.
@@ 3.2. Overview of WildGraph
  • WildGraph accepts a sequence of training samples of known trajectories, each of which consist of $m$ points. All of these points are used in the Hierarchical Network Generator, which implements a prototype graph ($\mathcal{G}$) based on them. A node represents a region of $\mathcal{G}$, while the edges represent transitions from one node to another as observed in a given trajectory.
  • A pre-training of the nodes of $\mathcal{G}$ is performed based on the connections identified. We utilize node2vec to generate representations of each node based on the input and output connections. The pre-training allows us to encode the relationships between the nodes and provides a representation learning of the transitions in the network, as well reduces the dimensions of the embeddings, compared to structures such as Bag of words (BOW).
  • The Variational Recurrent Network (VRN) aims to learn the empirical probability distribution of all states, i.e. learning the transition kernel, in order to generate trajectories following this distribution in an auto-regressive manner. Post-training, the output distribution of each node is aggregated in a latent dictionary (described below) that is used later during the generation process.
Once training is completed, the process of generating new trajectories utilizes the latent dictionary to stochastically select the next region based on the current region. The latent dictionary is a multi-valued dictionary that collects latent values ($z$) that correspond to a given state. The generation process is repeated until the desired sequence length (input parameter) is reached. For each region, an occupancy based sampling converts the region into a point. The occupancy sampler divides the entire area $\mathcal{A}$ into smaller regions, i.e. dots, and builds a discrete heatmap on the entire area. The process of converting a region to a point involves sampling a dot (a small region) from the corresponding region based on the heatmap, and randomly sampling a point from the dot.

### v2

# WildGraph: Realistic Long-Horizon Trajectory Generation with Limited Sample Size

@@ Abstract.
Trajectory generation is an important task in movement studies. Generated trajectories augment the training corpus of deep learning applications, facilitate experimental and theoretical research, and mitigate the privacy concerns associated with real trajectories. This is especially significant in the wildlife domain, where trajectories are scarce due to the ethical and technical constraints of the collection process. In this paper, we consider the problem of generating long-horizon trajectories, akin to wildlife migration, based on a small set of real samples. We propose a hierarchical approach to learn the global movement characteristics of the real dataset, and recursively refine localized regions. Our solution, WildGraph discretizes the geographic path into a prototype network of H3111https://www.uber.com/blog/h3/ regions and leverages a novel recurrent VAE to probabilistically generate paths over the regions, based on occupancy. Experiments performed on two wildlife migration datasets demonstrate the remarkable capability of WildGraph to generate realistic months-long trajectories using a sample size as small as 60 while improving generalization compared to existing work. Moreover, WildGraph achieves superior or comparable performance on performance measures, including geographic imagery similarity. Our code is published on the following repository: https://github.com/aliwister/wildgraph.
@@
journalyear: 2024conference: ACM Conference; 2024; University Park, PA, USAccs: Computing methodologies Model development and analysisccs: Computing methodologies Machine learning approachesccs: Computing methodologies Simulation toolsccs: Information systems Geographic information systems
@@ 1. Introduction
Recent advances in wildlife tracking have played a key role in contributing spatial trajectory datasets to wildlife researchers. These datasets provide invaluable insights into the patterns and behaviors of movement, enabling a deeper understanding of the underlying mechanisms driving wildlife movement. In particular, wildlife trajectories are utilized by scientists across different fields to promote wildlife conservation (Wall et al., 2014), anti-poaching efforts (Park et al., 2015), mitigating human-wildlife conflict (Buchholtz et al., 2020), and to help guide policy towards wildlife preservation (Tomkiewicz et al., 2010).
However, most published datasets are scarce, i.e. they include a small sample set of animals. Wildlife researchers have cited ethical and technical constraints involved in wildlife trajectory data collection (Dore et al., 2020). In particular, data is traditionally collected using devices attached directly to the subject, which requires capturing and possibly sedating it. The devices are further constrained by the animal’s size, and may be prone to intermittent or permanent failures, thus limiting the availability of trajectory data.
To overcome the challenges associated with collecting real movement data, trajectory generation has been considered in various domains, including human mobility (Long et al., 2023; Yuan et al., 2022; Rossi et al., 2021b), vehicle mobility (Cao and Li, 2021; Rossi et al., 2021a), as well as wildlife (Al-Lawati et al., 2023; Technitis et al., 2015) movement studies. Generated trajectories help enrich the training corpus to advance deep learning (DL) models (Castelli, 2023), and identify trends and patterns. For example, with trajectory augmentation (Zhou et al., 2021; J. Haranwala et al., 2023), known past poaching incidents mapped against a large trajectory set ($real$cup $generated$) help enforcement efforts to target high-risk regions (Xu et al., 2020). Furthermore, trajectory generation plays a key role in addressing privacy concerns of releasing real data, which may put wildlife at the risk of poaching (Choudhary et al., 2023). Generated trajectories also provide valuable insight into experimental studies, scenario planning, and risk assessment. Researchers can simulate the effects of policies and environmental decisions, thereby improving the robustness and generalization of their models and analyses (Beach and Clark, 2015; Catano et al., 2015).
Generative models powered by deep learning algorithms have demonstrated impressive capabilities in learning the knowledge semantics of a dataset in various domains and capturing it to produce original data based on the learned patterns. Simple methods based on variational autoencoders (VAEs) have shown a tremendous capability to generate meaningful samples based on a small data sample, i.e. the small data problem. Ideally, we want the generated set of trajectories to have the same characteristics (measured by, e.g. Hausdorff Distance or correlation coefficient) of a set of real trajectories, and generates original samples that preserve the patterns of the real trajectories (e.g., clusters or distance similarity). For example, in Figure 1, it can be visually verified that WildGraph generates much better trajectories, than using a Generative Adversarial Network (GAN) for this problem.
@@
However, existing solutions in human and vehicle mobility fail to generalize to the wildlife domain. In particular, wildlife trajectory generation presents a unique set of challenges. First is the small data problem: real trajectory datasets are sparse and include a small number of subjects. Second, the length of each trajectory tends to span several months, or years (e.g. in that long timespan, they may vary due to confounding factors, weather patterns, extreme events that are outliers, etc.) to provide practical value. Finally, often there is no established network information, e.g. roads, which wildlife may follow, (although in some land animals, there may be tracks on the ground trailed by the migrations, these are not rigid and animals often stray from them to create new trails/paths easily) thus increasing the complexity of the generation task.
Despite these challenges, interestingly, simple deep learning-based techniques, such as VAEs have demonstrated promising results in generating long-horizon trajectories using a small set of real training samples (Chen et al., 2021; Al-Lawati et al., 2023). VAEs accomplish this by mapping trajectories in a probabilistic manner to a latent space and teaching the decoder to reconstruct the encoded trajectory from the latent representation. Hence, new trajectories that have similar properties as the real trajectory may be generated by sampling the latent space.
@@
In this paper, we address these challenges by constructing a prototype network of regions for wildlife movement that models trajectories as a path across the network. We utilize the set of known trajectories to construct the global regions, and recursively refine localized regions, aka nodes, into multiple granular nodes. This helps preserve the global direction of movement, and at the same time captures the local properties of the generated trajectories. By adjusting the region sizes of the network to restrict the selection of potential points from a particular region, we improve the generalization of the generated trajectories, and provide a better outcome across multiple benchmarks.
@@
Once each known trajectory is encoded as a path over the network, we process it using a Variational Recurrent Network (VRN), which trains a model that generates a dictionary of latent spaces for each region, as depicted in Figure 3. The details of our framework are discussed in further details in Section 3.
@@
  • We propose WildGraph, a framework that generates realistic long-horizon wildlife trajectories based on a sparse set of real trajectories using a novel VRN implementation.
@@
  • We conduct multiple experiments against two wildlife datasets, and benchmark our methods against our earlier work, WildGEN (Al-Lawati et al., 2023), a VAE based approach, a GAN based approach, and a Transformers-based approach based on the prototype network. We use various comparison measures including a geographic imagery-based measure using SatCLIP embeddings (Klemmer et al., 2023) to evaluate the geographic similarity between trajectories.
To the best of our knowledge, this paper along with our previous workshop paper: WildGEN (Al-Lawati et al., 2023) are the first to propose deep learning-based trajectory generation for wildlife movement. Specifically, WildGEN demonstrated that deep learning approaches produce significantly more realistic trajectories than traditional methods, such as Correlated Random Walk (CRW) (Renshaw and Henderson, 1981) and Levy Flight (zaburdaev2015levy). This study builds upon our previous work by refining the point selection process along trajectories, as assessed by the similarity between generated trajectories and real trajectories in our test set.
As a result of the lack of recent application-specific benchmarks, and the key differences of trajectory generation applications in other mobility areas (e.g. based on a network), we compare our method with multiple standard generative algorithms, including VAE, Transformer, and GANs.
Coarser zoom level
@@ 2. Related Work
In this section, we review recent works from different domains related to the methods and solutions presented in our paper.
Trajectory Generation. In the pre-deep-learning era, trajectory generation was relatively subject specific. In ecology, researchers used methods such as Levy Flight (zaburdaev2015levy), Correlated Random Walk (CRW) (Technitis et al., 2015), and other velocity based methods (Technitis et al., 2015; Ahearn et al., 2017) that rely predominantly on the subject (animal) considered. The advent of deep-learning-based generative models such as VAEs and GANs resulted in some of the earliest attempts at using these approaches for trajectory generation. Chen et al. (Chen et al., 2021) benchmarked GANs and VAEs for the task of generation, and found that VAEs provide more consistent results. Our earlier work, WildGEN, was among the first methods to address the problem of deep-learning-based trajectory generation for wildlife (Al-Lawati et al., 2023). We utilized VAEs with added layers of post-processing that mitigated some of the challenges discussed above. However, much like other VAE-based methods, it suffers from limited generalization.
@@
Trajectory Similarity. Trajectory similarity is a complementary problem to trajectory generation since the goal of generation is to produce a set of trajectories similar to a set of known trajectories. An interesting approach to the problem was considered by Yao, et al. (Yao et al., 2022), where a region is recursively subdivided to capture the region-based overlap similarity between compared trajectories. Similar to our approach, they recursively split regions into sub-regions to create a balanced graph. Next, they utilize a graph-based attention layer to model the embedding of each node. In comparison, our approach utilizes node2vec (Grover and Leskovec, 2016) to overcome the over-parameterization in our problem given the shape of the training data. Furthermore, they represent the trajectory using an adjacency matrix which is not permutation invariant. A time-aware approach to generate similar graph snapshots at each step for a trajectory generation problem would render the approach considerably inefficient. Moreover, it will fail in the presence of cycles in the graph, as it cannot correctly determine the sequence of nodes in the trajectory.
@@
Graph Generation. Early generative graph models were based on adjacency matrix generation using encoder-decoder structures (You et al., 2018; Simonovsky and Komodakis, 2018). However, these approaches fail to capture the dynamic nature of spatio-temporal graphs. Recently, models such as STGD-VAEs (Du et al., 2022), and STGEN (Ling et al., 2023) were proposed to address these limitation. STGD-VAEs factor graphs into dependent ordinal values and jointly optimize the objective. On the other hand, STGEN utilizes the distribution of spatiotemporal walks to adversarially learn the characteristics of the graph. However, multi-objective approaches appear to be a poor fit for this problem, given the small sample sizes, and long horizon. We validate this in our experiments and demonstrate the GANs provide poor results on all metrics (see Fig 1).
@@
Sequence Modeling. Sequence-to-Sequence models have been used for various tasks in wildlife trajectory studies, including prediction (Feng and Gordon, 2019), imputation (Li et al., 2021) and interpolation (Wan and Dodge, 2023). They identify implicit patterns by using LSTM or RNN memory trained with new observations in an auto-regressive manner (Park et al., 2018). However, the multivariate nature of spatial data often results in a quick deterioration of predictions on tasks involving long horizon prediction (Sagheer and Kotb, 2019), particularly in small data settings. To address this, Li et al. (Li et al., 2021) attempted to simplify the problem to a prediction of step size instead of location, but this approach offers very limited usefulness. Our work adopts a more practical approach by leveraging the properties of the H3 spatial library to represent geospatial regions, where each region is encoded with a string representing its address encoded within the library, thus reducing the problem of generating a path to one in a single dimension.
@@
In addition, encoding each region’s connection (i.e., the probability of whether a bird will fly from one region to another) with other regions provides additional context to the problem, and facilitates generating long-horizon trajectories based on a small set of training samples.
@@ 3. Preliminaries
A good generated trajectory accurately reflects the natural movement patterns and behaviors of the wildlife. A key measure is how well the model can predict trajectories it has not encountered, i.e., a held-out set.
@@ 3.1. Problem Statement
Formally, given a set of $n$ real trajectories in an area $\mathcal{A}$, each consisting of $m$ points, the objective is to produce a set of synthetic trajectories that maximizes the following:
@@
    (Path Similarity) similarity between the points of the synthetic trajectories to that of the real trajectories.
@@
  3. (3)
    (Coverage) The percentage of samples in the held-out set closest to any generated sample, where “closest” is defined using a distance metric (we leverage our path similarity metrics). This could be interpreted as a measure of how well the generated samples represent the true distribution of the held-out set. This measure is particularly critical in our context, as it emphasizes the feasibility and representativeness of the generated trajectories.
The task is to learn a model to generate synthetic trajectories that maximize the similarity between a test set comprising of $k$ real trajectories $T_{test}=T_{1},T_{2},\ldots,T_{k}$, each comprising $m$ spatio-temporal points, where $T_{i}=(p_{1}^{i},p_{2}^{i},\ldots,p_{m}^{i})$, with $p_{j}^{i}$ representing latitude ($y$) and longitude ($x$) values. The objective is to find a corresponding set $S={S_{1},S_{2},\ldots,S_{t}}$ of synthetic trajectories that best matches the (test set of) real trajectories.
To achieve this, we aim to maximize the overall similarity between $T_{test}$ and $S$, denoted as:
$\max\text{ Similarity}(T_{test},S)$
Here, the similarity measure encompasses both Path similarity, Likeness, and Coverage. Note that using a held-out set to compare with the generated trajectories, post-training, ensures robustness and generalization of the proposed solution. Further details on the evaluation metrics utilized are provided in Section 3.5.
@@ 3.2. Overview of WildGraph
  • WildGraph accepts a sequence of training samples of known trajectories, each of which consists of $m$ points. All of these points are used in the Hierarchical Network Generator, which implements a prototype graph ($\mathcal{G}$) based on them. A node represents a region of $\mathcal{G}$, while the edges represent transitions from one node to another as observed in a given trajectory.
  • A pre-training of the nodes of $\mathcal{G}$ is performed based on the connections identified. We utilize node2vec to generate representations of each node based on the input and output connections. The pre-training allows us to encode the relationships between the nodes and provides a representation learning of the transitions in the network, as well as reduces the dimensions of the embeddings, compared to structures such as Bag of words (BoW).
  • The VRN aims to learn the empirical probability distribution of all states, i.e. learning the transition kernel, in order to generate trajectories following this distribution in an auto-regressive manner. Post-training, the output distribution of each node is aggregated in a latent dictionary (described below) that is used later during the generation process.
Once training is completed, the process of generating new trajectories utilizes the latent dictionary to stochastically select the next region based on the current region. The latent dictionary is a multi-valued dictionary that collects latent values ($z$) that correspond to a given state. The generation process is repeated until the desired sequence length (input parameter) is reached.
For each region, an occupancy based sampling converts the region into a point. The occupancy sampler divides the entire area $\mathcal{A}$ into smaller regions, i.e. dots, and builds a discrete heatmap on the entire area. The process of converting a region to a point involves sampling a dot (a small region) from the corresponding region based on the heatmap, and randomly sampling a point from the dot.